What atomic terms (including J values) are possible for the configuration np1nf1? what is the ground state term?
Answer #1

ANSWER:-

term symbol formula is 2S+1LJ

atomic term symbol for np1

  • L value for p +1 because there is only one electron so L = p because 0,1,2,3 stands for s,p,d,f...
  • 2S+1 value is ,

s= number of unpaired electrons present so there is only one electron

s = 1/2 x no of unpaired electrons

s = 1/2x 1 = 1/2

  • so 2S+1 = 2
  • ang J=[L-S] for less then half-filled subshell
  • so J = 1
  • so term symbol for np1 is 2P1

and for nf1 L value for f orbital should be 3

and 2S+1 = 2

and J=[L-S]

J = 1

so term symbol for nf1 is 2F1